OpenStudy (anonymous):

a circle has a circumference of 22 inches. what is the radius of the circle

OpenStudy (mayankdevnani):

C=pi*d

hartnn (hartnn):

circumference = 2pi r = 22 find r from here

OpenStudy (mayankdevnani):

3.14*2*r=22

OpenStudy (wwe123):

2r*pi =22 r=22/2 * 7/22 r=11*7/22 r=7/2